English
Tamil Nadu Board of Secondary EducationHSC Science Class 11

A problem in Mathematics is given to three students whose chances of solving it are 13,14 and 15. What is the probability that exactly one of them will solve it?

Advertisements
Advertisements

Question

A problem in Mathematics is given to three students whose chances of solving it are `1/3, 1/4` and `1/5`. What is the probability that exactly one of them will solve it?

Sum
Advertisements

Solution

`"P"(bar"A"_1) = 1 - "P"("A"_1) = 1 - 1/3 = 2/3`

`"P"(bar"A"_1) = 1 - "P"("A"_2) = 1 - 1/4 = 3/4`

`"P"(bar"A"_1) = 1 - "P"("A"_3) = 1 - 1/5 = 4/5`

Probability of Exactly one student solving the problem

= `"P" [("A"_1 ∩ bar"A"_2 ∩ bar"A"_3) ∪ (bar"A"_1 ∩ "A"_2 ∩ bar"A"_3) ∪ (bar"A"_1 ∩ bar"A"_2 ∩ "A"_3)]`

= `"P"("A"_1 ∩ bar"A"_2 ∩ bar"A"_3) + "P"(bar"A"_1 ∩ "A"_2 ∩ bar"A"_3) + "P"(bar"A"_1 ∩ bar"A"_2 ∩ "A"_3)`

= `"P"("A"_1) "P"(bar"A"_2) "P"(bar"A"_3) + "P"(bar"A"_1) "P"("A"_2) "P"(bar"A"_3) + "P"(bar"A"_1) "P"(bar"A"_2) "P"("A"_3)`

= `1/3 xx 3/4 xx 4/5 + 2/3 xx 1/4 xx 4/5 + 2/3 xx 3/ xx 1/5`

= `(12 + 8 + 6)/60`

= `26/16`

= `13/30`

[Probability of exactly one student solving the problem = Probability [(A1 solving the problem and A2, A3 non solving the problem) or (A1, A3 non solving and A2 solving) or (A1, A2 solving and A3 non solving)]

shaalaa.com
  Is there an error in this question or solution?
Chapter 12: Introduction to probability theory - Exercise 12.3 [Page 259]

APPEARS IN

Samacheer Kalvi Mathematics - Volume 1 and 2 [English] Class 11 TN Board
Chapter 12 Introduction to probability theory
Exercise 12.3 | Q 6. (ii) | Page 259

RELATED QUESTIONS

Assume that the chances of a patient having a heart attack is 40%. Assuming that a meditation and yoga course reduces the risk of heart attack by 30% and prescription of certain drug reduces its chance by 25%. At a time a patient can choose any one of the two options with equal probabilities. It is given that after going through one of the two options, the patient selected at random suffers a heart attack. Find the probability that the patient followed a course of meditation and yoga. Interpret the result and state which of the above stated methods is more beneficial for the patient.


Suppose that 80% of all families own a television set. If 5 families are interviewed at  random, find the probability that
a. three families own a television set.
b. at least two families own a television set.


A black and a red dice are rolled. 

Find the conditional probability of obtaining a sum greater than 9, given that the black die resulted in a 5.


A fair die is rolled. Consider events E = {1, 3, 5}, F = {2, 3} and G = {2, 3, 4, 5} Find P (E|G) and P (G|E)


Suppose we have four boxes. A, B, C and D containing coloured marbles as given below:

Box Marble colour
  Red White Black
A 1 6 3
B 6 2 2
C 8 1 1
D 0 6 4

One of the boxes has been selected at random and a single marble is drawn from it. If the marble is red, what is the probability that it was drawn from box A?, box B?, box C?


An urn contains 2 white and 2 black balls. A ball is drawn at random. If it is white, it is not replaced into the urn. Otherwise, it is replaced with another ball of the same colour. The process is repeated. Find the probability that the third ball is drawn is black.


A box has 20 pens of which 2 are defective. Calculate the probability that out of 5 pens drawn one by one with replacement, at most 2 are defective.


Three cards are drawn at random (without replacement) from a well-shuffled pack of 52 playing cards. Find the probability distribution of the number of red cards. Hence, find the mean of the distribution.


If events A and B are independent, such that `P(A)= 3/5`,  `P(B)=2/3` 'find P(A ∪ B).


In an examination, 30% of students have failed in subject I, 20% of the students have failed in subject II and 10% have failed in both subject I and subject II. A student is selected at random, what is the probability that the student has failed in subject I, if it is known that he is failed in subject II?


Two cards are drawn one after the other from a pack of 52 cards without replacement. What is the probability that both the cards drawn are face cards?


Select the correct option from the given alternatives :

Bag I contains 3 red and 4 black balls while another Bag II contains 5 red and 6 black balls. One ball is drawn at random from one of the bags and it is found to be red. The probability that it was drawn from Bag II


Can two events be mutually exclusive and independent simultaneously?


The probability that a car being filled with petrol will also need an oil change is 0.30; the probability that it needs a new oil filter is 0.40; and the probability that both the oil and filter need changing is 0.15. If a new oil filter is needed, what is the probability that the oil has to be changed?


A year is selected at random. What is the probability that it contains 53 Sundays


Suppose the chances of hitting a target by a person X is 3 times in 4 shots, by Y is 4 times in 5 shots, and by Z is 2 times in 3 shots. They fire simultaneously exactly one time. What is the probability that the target is damaged by exactly 2 hits?


If X denotes the number of ones in five consecutive throws of a dice, then P(X = 4) is ______ 


A bag contains 6 red and 5 blue balls and another bag contains 5 red and 8 blue balls. A ball is drawn from the first bag and without noticing its colour is placed in the second bag. If a ball is drawn from the second bag, then find the probability that the drawn ball is red in colour.


Let A, B be two events such that the probability of A is `3/10` and conditional probability of A given B is `1/2`. The probability that exactly one of the events A or B happen equals.


A Problem in Mathematics is given to the three students A, B and C. Their chances of solving the problem are `1/2, 1/3` and `1/4` respectively. Find the probability that at least two of them will solve the problem.


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×